OpenStudy (anonymous):

express as a single fraction : ((sqrt2)/2) x ((sqrt3)/2) + ((sqrt2)/2) x (1/2)

OpenStudy (anonymous):

\[= (\sqrt{6} / 4) + \sqrt{2}/4) = (\sqrt{6}+\sqrt{2}) \div 4\]